# Accelerating Molecular Dynamics Simulations with Foundation Neural Network Models using Multiple Time-Step and Distillation

### v2

@@ 3.3 Protein-ligand complexes
To validate our scheme on a biologically relevant system, we performed a $20\text{\,}\mathrm{n}\mathrm{s}$ NVT molecular dynamics simulation of the lysozyme–phenol complex (PDB ID: 4I7L) in explicit water. We first resorted to a setup of $4\text{\,}\mathrm{f}\mathrm{s}$ outer time step and HMR which is stable for bulk water, but this led to instabilities after $2\text{\,}\mathrm{n}\mathrm{s}$ of simulation. Figure 4(a) shows the velocity autocorrelation spectra computed over a short simulation with both STS and MTS with HMR enabled. The close similarity of the obtained spectra indicates that the instabilities are not due to the resonance phenomena described above.
@@
In practice, this source of instability lies in the presence of ”holes” in the potential energy surface of the small model, yielding large differences between its forces and the target ones. This is illustrated in Figure 4(b) showing the distribution of the norm of the difference of forces between FeNNix-Bio1(M) and the distilled model (either finetuned on system-specific fragments or the generic one) along an STS trajectory of the protein-ligand system. We observe a dense distribution of values close to $0\text{\,}\mathrm{k}\mathrm{c}\mathrm{a}\mathrm{l}\mathrm{/}\mathrm{m}\mathrm{o}\mathrm{l}\mathrm{/}\mathrm{\text{Å}}$ and a long tailed distribution starting from $150\text{\,}\mathrm{k}\mathrm{c}\mathrm{a}\mathrm{l}\mathrm{/}\mathrm{m}\mathrm{o}\mathrm{l}\mathrm{/}\mathrm{\text{Å}}$. The latter is associated to large nonphysical force discrepancies yielding the infrequent instabilities that we observe in the MTS simulation.
We attempted to mitigate this issue by clipping the force difference to exclude unrealistic values while preserving the meaningful ones (in practice, we used a threshold of $150\text{\,}\mathrm{k}\mathrm{c}\mathrm{a}\mathrm{l}\mathrm{/}\mathrm{m}\mathrm{o}\mathrm{l}\mathrm{/}\mathrm{\text{Å}}$ on force components). While improving the stability, this setup still leads to the rare occurrence of crashes due to large differences when the time step is equal to or larger than $4\text{\,}\mathrm{f}\mathrm{s}$. We expect that an active learning loop where the ”holes” are first identified and then included in a fine-tuning stage will systematically improve the stability of our MTS scheme with larger time steps; this will be the focus of a future work.
@@
b) Distribution (in log-scale) of the norm of the force differences for the generic model, the fine-tuned model trained on forces, energies, and atomic energies, and the fine-tuned model trained on forces and energy, with respect to the FENNIX-Bio1(M) reference model. The distributions are estimated over the first 500 frames of a phenol–lysozyme in water simulation.
Using an inner time step of $1.75\text{\,}\mathrm{f}\mathrm{s}$ and an outer time step of $3.5\text{\,}\mathrm{f}\mathrm{s}$ with HMR, we obtain stable simulations over $20\text{\,}\mathrm{n}\mathrm{s}$ while preserving the protein’s structure and the ligand’s binding mode as shown in Figure 5. This setup provides a speedup of 2.7 with respect to the STS simulation with $1\text{\,}\mathrm{f}\mathrm{s}$ time step, strongly accelerating production simulations with biological systems.
Figure 5: Time evolution of the protein backbone RMSD and the ligand’s Distance to Bound Configuration (DBC)46 during a $20\text{\,}\mathrm{n}\mathrm{s}$ simulation of the lysozyme–phenol complex in water. Results obtained with the implemented MTS integrator (solid lines) are compared to those from a reference STS simulation (dotted lines). MTS simulations use either the generic or finetuned models combined with the FeNNix-Bio1(M) potential with an internal time step of $1.75\text{\,}\mathrm{f}\mathrm{s}$ and an external time step of $3.5\text{\,}\mathrm{f}\mathrm{s}$ with HMR.
@@ 4 Conclusion
We have introduced a practical and efficient Multiple Time Step (MTS) scheme tailored to machine-learned force fields, enabling significant acceleration of molecular dynamics simulations without compromising accuracy for both static and dynamical properties. Our approach, based on a RESPA-like scheme, couples a fast, distilled neural network model with the accurate FeNNix-Bio1(M) reference potential. We demonstrated two complementary distillation strategies: (i) on-the-fly system-specific model distillation, and (ii) the use of a lightweight transferable model for broader applicability.
We achieve substantial speedups – up to 4 in mostly homogeneous systems (bulk water and solvated small molecules) and 2.7 in large protein-ligand complexes – while preserving key physical observables such as radial distribution functions, hydration free energies, diffusion coefficients, and protein-ligand structural properties. Note that this is a first estimation of the computational gains as the code has yet to be optimized further to handle the dual-level approach more efficiently. All in all, our approach allows to reach nearly $7\text{\,}\mathrm{n}\mathrm{s}$ of simulation per day on a single A100 GPU for a realistic protein-ligand complex while preserving the ab initio-like accuracy of the FeNNix-Bio1(M) model. In combination with well-established accelerated sampling schemes 47, 48, 49, this enables truly large-scale simulations with foundation neural neural network potentials.
Future work will focus on expanding the applicability of this approach through two main avenues. First, randomized time stepping could be explored to further increase the effective time step, while allowing for controlled velocity rescaling as proposed in stochastic RESPA variants, i.e. JUMP integrators 44. Second, system-specific models can be fine-tuned more efficiently using fragment-based active-learning strategies to progressively fill-out ”holes” in the small model’s potential energy surface, potentially improving stability with even larger time steps for more complex chemical systems.

### v3

@@ 3.3 Protein-ligand complexes
To validate our scheme on a biologically relevant system, we performed a $20\text{\,}\mathrm{n}\mathrm{s}$ NVT molecular dynamics simulation of the lysozyme–phenol complex (PDB ID: 4I7L) in explicit water. Because of the chemical diversity of this system, we do not train a system-specific model from scratch. We thus first assess the ability of the generic model to handle the system and then show that its performance can be improved via active-learning-driven fine tuning.
@@
Using an inner time step of $1.75\text{\,}\mathrm{f}\mathrm{s}$ and an outer time step of $3.5\text{\,}\mathrm{f}\mathrm{s}$ with HMR, we obtain stable simulations over $20\text{\,}\mathrm{n}\mathrm{s}$ with the generic model while preserving the protein’s structure and the ligand’s binding mode as shown in Figure 5. Note that simulations with a $4\text{\,}\mathrm{f}\mathrm{s}$ outer time step and HMR (which were stable for homogeneous systems) displayed instabilities after $2\text{\,}\mathrm{n}\mathrm{s}$. The close similarity of the velocity autocorrelation spectra (Figure 4(a)) between STS and DMTS indicates that these instabilities are not due to the resonance phenomena described above. They instead originate from the presence of “holes” in the potential energy surface of the small model, yielding large differences between its forces and the target ones. This is illustrated in Figure 4(b) showing the distribution of the norm of the difference of forces between FeNNix-Bio1(M) and the distilled model along an STS trajectory of the protein-ligand system. We observe a dense distribution of values close to $0\text{\,}\mathrm{k}\mathrm{c}\mathrm{a}\mathrm{l}\mathrm{/}\mathrm{m}\mathrm{o}\mathrm{l}\mathrm{/}\mathrm{\text{Å}}$ and a long-tailed distribution starting from $150\text{\,}\mathrm{k}\mathrm{c}\mathrm{a}\mathrm{l}\mathrm{/}\mathrm{m}\mathrm{o}\mathrm{l}\mathrm{/}\mathrm{\text{Å}}$. The latter is associated with large nonphysical force discrepancies yielding the infrequent instabilities that we observe in the DMTS simulation.
To address this limitation, we introduce two complementary strategies:
Small message-passing model. First, we explored enriching the small model architecture with explicit message passing. The intuition is that a more expressive model, with more parameters and deeper local information flow, should exhibit fewer “holes” in the potential energy surface, thereby reducing the occurrence of pathological configurations. This strategy proved effective: with a single message passing layer, we obtained stable $20\text{\,}\mathrm{n}\mathrm{s}$ simulations using $1\text{\,}\mathrm{f}\mathrm{s}$-$4\text{\,}\mathrm{f}\mathrm{s}$ inner-outer time steps, confirming the improved robustness of the small model (see Figure S2 in Supporting information). However, the added architectural complexity significantly increases the inference cost, making the approach less attractive for large-scale production simulations (going from $6.54\text{\,}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{d}\mathrm{a}\mathrm{y}$ for the generic model to $5.95\text{\,}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{d}\mathrm{a}\mathrm{y}$ for the message passing model).
Improving the small model via active learning. We therefore investigated an alternative strategy that preserves the simplicity of the original model: active-learning-driven fine tuning. In this approach, we design an active-learning procedure that automatically detects frames and atoms where the force difference exceeds an unrealistic threshold of $150\text{\,}\mathrm{k}\mathrm{c}\mathrm{a}\mathrm{l}\mathrm{/}\mathrm{m}\mathrm{o}\mathrm{l}\mathrm{/}\mathrm{\text{Å}}$. For such frames, the integrator temporary reverts to the STS integrator and local clusters centered on the problematic atoms are added to an adaptive fine-tuning dataset. In total, curating the final dataset, required approximately $400\text{\,}\mathrm{p}\mathrm{s}$ of simulation (see details in Supporting Information). The generic small model is then refined on this curated dataset, systematically improving its robustness in regions of configuration space previously associated with instabilities.
Remarkably, this refinement procedure achieved, without increasing the model’s complexity, stable $20\text{\,}\mathrm{n}\mathrm{s}$ simulations using up to $2\text{\,}\mathrm{f}\mathrm{s}$-$4\text{\,}\mathrm{f}\mathrm{s}$ inner-outer time step. The protein’s structure and the ligand’s binding mode were preserved as shown in Figure 5. Furthermore, we show in the Supporting Information (Figure S1) that the potential energy is stable along the whole $20\text{\,}\mathrm{n}\mathrm{s}$ simulations in all cases. With this configuration, we reach a production speed of $7.45\text{\,}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{d}\mathrm{a}\mathrm{y}$, corresponding to a $2.92\text{\,}$ speed up compared to the STS simulations, strongly accelerating production simulations with biological systems. This demonstrates that targeted fine tuning can rival the stability gains of more expressive architectures while retaining a significantly lower inference cost. Looking ahead, this adaptive strategy could be further enhanced through more sophisticated active-learning schemesvanderoordHyperactiveLearningDatadriven2023, kulichenkoUncertaintydrivenDynamicsActive2023.
@@
b) Distribution (in log-scale) of the norm of the force differences for the generic model with respect to the FENNIX-Bio1(M) reference model. The distributions are estimated over the first 500 frames of a phenol–lysozyme in water simulation.
Figure 5: Time evolution of the protein backbone RMSD and the ligand’s Distance to Bound Configuration (DBC)salari2018streamlined during a $20\text{\,}\mathrm{n}\mathrm{s}$ simulation of the lysozyme–phenol complex in water. Results obtained with the implemented MTS integrator (solid lines) are compared to those from a reference STS simulation (dotted lines). DMTS simulations use either the generic or active learned models combined with the FeNNix-Bio1(M) potential with an internal time step of $1.75\text{\,}\mathrm{f}\mathrm{s}$ and an external time step of $3.5\text{\,}\mathrm{f}\mathrm{s}$ with HMR for generic model and an internal-external time step of $2\text{\,}\mathrm{f}\mathrm{s}$-$4\text{\,}\mathrm{f}\mathrm{s}$ with HMR for the active learned model.
@@ 4 Conclusion
We have introduced a practical and efficient Distilled Multiple Time Step (DMTS) scheme tailored to machine-learned force fields, enabling significant acceleration of molecular dynamics simulations without compromising accuracy for both static and dynamical properties. Our approach, based on a RESPA-like scheme, couples a fast, distilled neural network model with the accurate FeNNix-Bio1(M) reference potential. We demonstrated two complementary distillation strategies: (i) on-the-fly system-specific model distillation, and (ii) the use of a lightweight transferable model for broader applicability.
We achieve substantial speedups – up to 4 in mostly homogeneous systems (bulk water and solvated small molecules) and 2.92 in large protein-ligand complexes – while preserving key physical observables such as radial distribution functions, hydration free energies, diffusion coefficients, and protein-ligand structural properties. Note that this is a first estimation of the computational gains as the code has yet to be optimized further to handle the dual-level approach more efficiently. All in all, our approach allows to reach above $7\text{\,}\mathrm{n}\mathrm{s}$ of simulation per day on a single A100 GPU for a realistic protein-ligand complex while preserving the ab initio-like accuracy of the FeNNix-Bio1(M) model. In combination with well-established accelerated sampling schemes Celerse2022, D1SC00145K, ansari2025lambda, this enables truly large-scale simulations with foundation neural network potentials.
Future work will focus on expanding the applicability of this approach through two main avenues. First, randomized time stepping could be explored to further increase the effective time step, while allowing for controlled velocity rescaling as proposed in stochastic RESPA variants, i.e. JUMP integrators velocity_jumps. Second, combining the fragment-based active-learning strategy with the generic message passing architecture offer a promising route forward. An active learning driven message passing model could selectively refine unstable regions on the fly, potentially enabling inner time step beyond $4\text{\,}\mathrm{f}\mathrm{s}$ while offsetting the higher inference cost of the message passing.